7. Find the area of a circle with a circumference of 31.4 units.
Answer by checkley77(12569)

Circumferance=2pir
31.4=2*3.14*r
31.4=6.28r
r=31.4/6.28
r=5 units is the radius.
area=pir^2
area=3.14^5^2
area=3.14*25
area=78.5 units for the area.